# Optimization under Attack: Resilience, Vulnerability, and the Path to Collapse

### v1

@@ III-B Network Model
Consider a network with $\mathsf{n}$ agents, denoted by $A$, each identified by a unique ID in the set $\{1,2,3,...,\mathsf{n}\}$. The network topology is represented as a connected graph $G=(A,E)$, where $A$ is the set of agents, and $E$ is the set of edges, such that $(j,i)\in E\neq(i,j)\in E$. Agents interact within a self-organized network through bidirectional communication to exchange information and update their states to align with the system goals.
Table II defines the notations used throughout the paper to formalize the network model and optimization framework.
TABLE II: Nomenclature utilized in this research
@@ III-C Optimization Framework in Discrete-Choice Scenarios
Agents balance their individual preferences with system-wide goals, which often involve conflicting criteria. Each agent has an individual preference for its plans, calculated by the discomfort cost ($D$), such that $D$=${}_{a,i}=f_{D}(p_{\mathsf{a,i}})$, where $f_{D}$ is the agent’s preference associated with a particular plan. Each agent $a$ measures the costs $D$ for each possible plan $p_{\mathsf{a,i}}$$\in$ $P_{\mathsf{a}}$ and the plan with the minimum cost is selected. While agents aim to minimize their own discomfort, they may also consider system-wide metrics such as the inefficiency cost ($I$). The inefficiency cost ($I$) is the measure used to evaluate the collective system-wide performance based on the aggregated responses of all agents. It represents the system-wide performance inefficiency that agents aim to minimize through coordinated decision-making: $I=f_{I}\left(\sum_{a=1}^{\mathsf{n}}(s_{\mathsf{a}})\right)$.
Agents optimize their decisions by minimizing a weighted combination of these costs, as illustrated in Equation 1. The behavior of agent $a$ is modeled by the corresponding weights $\beta_{a},\text{ and }\alpha_{a}$, prioritizing efficiency and comfort, respectively. A higher weight indicates a higher preference for minimizing the corresponding objective. On the other hand, a weight of 0 means that the corresponding objective is not considered. For example, When $\beta_{a}=0,\text{ and }\alpha_{a}=1$, the agent prioritizes system-wide efficiency, acting as altruistic. Conversely, when $\beta_{a}=1,\text{ and }\alpha_{a}=0$, the agent prioritizes its own interest, behaving as selfish.
@@
$=\arg\min_{i=1}^{k}\alpha_{a}\cdot f_{I}\left(s_{1},s_{2},\ldots,s_{\mathsf{n}}\right)$
$\quad+\beta_{a}\cdot f_{D}\left(D_{1,s},D_{2,s},\ldots,D_{\mathsf{n},s}\right)$ (1)
@@
The decision-making process in distributed optimization is influenced not only by each agent’s individual objectives but also by the collective impact of their choices on system-wide goals. This interdependence introduces significant complexity, particularly when cost functions are non-linear, as agents’ choices depend on one another. Minimizing inefficiency and discomfort costs under such conditions is an NP-hard problem [49]. These challenges are further exacerbated by the lack of complete information about other agents’ selections and the need for coordination among agents’ choices. Addressing these issues requires a decentralized optimization framework that enables agents to iteratively refine their plan selections based on aggregated contributions and prior iterations (see Section IV-A).
@@ III-D Adversarial Distributed Optimization Model
To study the resilience of distributed optimization in an adversarial environment, we partition the set of agents, $A$, into two subsets: adversary agents ($A_{\mathsf{d}}$) and legitimate agents ($A_{\mathsf{l}}$). Adversary agents, $A_{\mathsf{d}}\subseteq A$, prioritize their individual interests over collective system goals by adapting their behavior to maximize personal benefits. This behavior disregards system-wide efficiency and leads to suboptimal outcomes for the overall system. Legitimate agents, $A_{\mathsf{l}}\subseteq A$, align their actions with system-wide objectives to optimize overall performance. The total set of agents is defined as $A=A_{\mathsf{d}}\cup A_{\mathsf{l}}$.
Adversarial behavior amplifies the inefficiency cost $I$, reflecting the trade-off between minimizing individual discomfort ($D$) and optimizing overall system efficiency ($I$). While $D$ focuses on individual preferences, $I$ addresses the system-wide inefficiency caused by deviations from optimal resource allocation, underscoring the conflict between personal and collective goals. For instance, in a bike-sharing system, optimization ensures a balanced distribution of bikes across stations to meet user demand. Legitimate users select pick-up and drop-off stations with consideration for system-wide efficiency, aiming to maintain network equilibrium. Conversely, adversary users prioritize their convenience, choosing stations that minimize their discomfort without regard for the network state. This behavior creates imbalances, such as empty or overstocked stations, reducing overall efficiency and user satisfaction.
@@ IV Experimental Methodology
IV-A Decentralized Optimization Framework
The proposed model is implemented within the Iterative Economic Planning and Optimized Selections (I-EPOS) [^2] Available at https://github.com/epournaras/EPOS, a collective learning framework. I-EPOS employs a self-organized, multi-level hierarchical network structure that ensures efficient communication and coordinated decision-making while reducing communication overhead [51].
I-EPOS enables the agents to iteratively adjust their plan selections based on the aggregated plans of the agents beneath and previous iteration choices. The proposed model operates iteratively, enabling agents to coordinate their choices in collective decision-making. Each iteration consists of two distinct phases: a bottom-up phase and a top-down phase. During the bottom-up phase, agents select plans based on the aggregated choices of agents in the subtree beneath, as well as the selections made by all agents in the previous iteration. Conversely, the top-down phase addresses incomplete knowledge from higher branches in the hierarchy, enabling agents to revert to previous selections if no cost reduction is achieved. This process continues until a predefined iteration limit is reached or no further improvement in the optimization objective occurs [49, 50].
The hierarchical network structure is organized as an acyclic graph, where each parent agent aggregates responses from its children without redundancy, explicitly avoiding double counting of contributions from lower nodes in the hierarchy. This design ensures efficient coordination and alignment of individual decisions with system-wide objectives while maintaining overall optimization [49, 51, 52]. I-EPOS is well-suited for adversarial scenarios in Smart City applications due to its scalability, adaptability to diverse agent behaviors, and ability to handle adversarial conditions [50, 53].
@@ IV-B Experimental Setup
Experiments are conducted using multiple HPC servers with varying configurations that support large-scale simulations and ensure computational efficiency. These include high-memory nodes (up to 768 GB) and multi-core processors (up to 40 cores per node). In addition to these servers, the University of Leeds ARC4 system  [^3] ARC4 is an HPC cluster at Leeds providing a Linux-based HPC service based on CentOS 7. More information: https://arcdocs.leeds.ac.uk/systems/arc4.html is utilized. The ARC4 system includes two nodes, each equipped with 40 cores, 768 GB of memory, and 800 GB of storage, providing robust computational capacity for large-scale experiments.
@@ IV-C Application Scenarios
TABLE III: Description of the datasets and experimental setup in this research
Dataset Name No. Agents No. Plans Plan Size Agents Discomfort Cost Inefficiency Cost Total Experiments Energy 1000 10 144 Households Agent preference based on time shift Energy demand 30,018,000 Data-collective 72 3 64 participants Privacy loss Data collective inefficiency (quality of collected data) 65,433,600 (2 target signals) Voting 266 31 5 Voters Compromise Polarization 97,551,840 (120 target signals)
@@ IV-C2 Privacy dataset
The privacy dataset originates from a living-lab experiment at the Decision Science Laboratory [^5] https://www.descil.ethz.ch of ETH Zurich, involving 72 participants evaluating 64 data-sharing scenarios that involve 4 sensor types, data collectors, and contexts [53]. The data-sharing choices of each participant determine three data-sharing plans, representing their intrinsic motivation to share and two rewarded scenarios. The plans are assessed using privacy valuation schemes assigning normalized costs ranging from 0 to 1, where lower costs indicate less privacy compromise. The dataset facilitates testing under high and low privacy-preservation target signals, reflecting intrinsic preferences for data-sharing behaviors. The inefficiency cost is calculated based on the reduction in the quality of the aggregated data due to adversarial participants prioritizing minimal data sharing over collective goals. Adversarial participants disrupt coordination by focusing solely on minimizing their data sharing, under-mining the quality of service of data collectors.
@@ IV-C3 Voting dataset
This new dataset is derived from voting data in a regional election with five candidates and 266 voters [^6] The UK Labour Party Leadership Vote Available at https://preflib.simonrey.fr/datasets [54]. Each voter has 31 alternative voting plans, representing ranked preferences among the five candidates. The optimization focuses on minimizing polarization in decision outcomes, which refers to the extent of divergence in voters’ preferences from a collectively agreed compromise. Inefficiency cost is measured as the cost of polarized decision outcomes. To ensure fair representation, 120 target signals are generated from all combinations of values 0, 0.25, 0.5, 0.75, and 1. Adversarial behavior occurs when voters prioritize minimizing their personal compromise cost, disregarding collective consensus, which amplifies polarization and disrupts the system ability to achieve shared, balanced outcomes.
Table III provides an overview of each dataset used in this research, including the number of agents, the number and size of plans per agent, agent representation, the corresponding discomfort and inefficiency cost functions, and the total number of experiments conducted.
@@ IV-D Varying Dimensions and Performed Experiments
The experiments systematically evaluate the resilience and vulnerability of the system across the following dimensions:
@@
  • Adversarial severity ($\beta$): The adversarial preference to minimize discomfort cost ($\beta$) is varied across 30 levels, with $\beta$ incrementing from 0 to 1 such as $\beta=\frac{b}{30}$ for $b=\{1,2,3,...,30\}$. This variation captures the impact of varying levels of adversarial intent, where agents prioritize their personal objectives at the expense of system-wide goals, on overall system performance.
  • Adversary position: The influence of adversary positions within the hierarchical network is analyzed using two approaches: layer-wise and cumulative structural analysis. A binary tree structure is employed, with each hierarchical layer containing approximately $\log_{2}|A|$ agents, where $|A|$ is the total number of agents. The layer-wise analysis evaluates inefficiency costs across individual hierarchical layers under varying adversary scales (25%, 50%, 75%, 100%) and behavior levels ($\beta$ = 0.1, 0.4, 0.7, 1.0). The cumulative analysis examines the aggregated impact of adversary agents propagating through the network in top-down (root-to-leaf) and bottom-up (leaf-to-root) configurations.
[…]
@@ IV-E Evaluation Metrics
Optimization objectives: Three metrics evaluate system performance: inefficiency cost $I$, discomfort cost $D$, and compromise cost of legitimate agents. The inefficiency cost measures the overall cost incurred by the system. The discomfort cost evaluates agents’ dissatisfaction with non-preferred choices. The compromise cost quantifies the extent to which legitimate agents adjust their individual preferences to align with collective goals under adversarial influence. This metric is calculated as the deviation between the selected plans of legitimate agents and their most preferred plans. These metrics evaluate the trade-offs between individual agent goals and system-wide objectives, particularly under varying adversarial conditions.
Pareto optimality: In multi-objective optimization, the Pareto front represents the set of non-dominated solutions where no objective can be improved without compromising another. The knee point on the Pareto front indicates the solution that achieves the best trade-off between competing objectives, making it a key decision-making criterion [55, 56].
This study identifies the knee point using the Minimum Manhattan Distance (MMD) method, which measures the distance between each solution on the Pareto front and an ideal reference point where both objectives are optimized. The solution with the smallest distance to the ideal point is selected as the knee point. This approach ensures a balanced evaluation of discomfort cost ($D$) and inefficiency cost ($I$) while maintaining alignment with established methodologies in the literature [56].
Resilience, vulnerability, and collapse framework: To define thresholds for resilience, vulnerability, and collapse in distributed optimization, we apply the multi-Otsu thresholding method. The Otsu method is an efficient technique for dividing data into distinct regions by minimizing intra-class variance [57, 58]. In this study, the method segments data into three regions based on inefficiency levels: resilience (low inefficiency), vulnerability (moderate inefficiency), and collapse (high inefficiency). This framework provides a robust method for segmenting inefficiency data into resilience, vulnerability, and collapse states, enabling a clear assessment of system performance transitions under adversarial conditions.
V Results Analysis and Discussion
This section presents the results of extensive experiments evaluating the impact of adversarial agents on multi-objective distributed optimization across the Energy, Voting, and Privacy datasets.
@@ V Results Analysis and Discussion
(c) Compromised discomfort costs for non-adversary agents
Figure 1: Multi-objective optimization performance across different numbers of adversary agents and various levels of adversaries within the Energy, Voting, and Privacy datasets. R, V, and C indicate regions of Resilience, Vulnerability, and Collapse.
V-A Resilience Analysis
Figure 1 illustrates inefficiency costs, discomfort costs, and compromised discomfort costs across varying scale of adversaries and behavior levels. These metrics collectively capture system-wide inefficiency, individual agent dissatisfaction, and the trade-offs faced by legitimate agents under adversarial conditions.
Energy Dataset
[…]
Voting Dataset
The Voting dataset demonstrates significant stability, maintaining resilience across more than 90% of the operational space. Inefficiency remains low for $\beta<0.9$ and scale of adversaries below 50%. Vulnerability emerges between 50% and 70% adversary agents, with inefficiency increasing by 111%. Collapse occurs at $\beta\geq 0.96$ and proportions exceeding 70%. Discomfort costs decline gradually in resilient regions and are entirely eliminated under collapse conditions, highlighting a clear trade-off between individual comfort and system-wide efficiency. Compromised discomfort costs show resilience for adversary agents below 20%, remaining low even under high levels of adversarial behavior. However, collapse is observed for proportions exceeding 50% and adversarial levels $\beta>0.7$, where compromised discomfort costs rise smoothly. This trend suggests that legitimate agents in the Voting dataset show lower resilience compared to the Energy dataset.
Privacy Dataset
The Privacy dataset reveals lower resilience thresholds compared to the Energy and Voting datasets, especially for high privacy-preserving signal. For $\beta<0.2$ and adversary agents below 20%, inefficiency costs remain low. However, vulnerability emerges as proportions increase to 50%, and collapse occurs at $\beta\geq 0.3$ for high data-sharing levels and $\beta\geq 0.5$ for low data-sharing levels when adversary agents exceed 50%.
Discomfort costs follow similar trends, with resilience observed below 30% adversary agents for high privacy-preserving signal and 20% for low privacy-preserving signal. Collapse occurs when $\beta>0.3$ and proportions exceed 60%. Compromised discomfort costs rise steadily, peaking at values of 9 and 10 for high and low privacy-preserving signal, respectively, at 90% adversary agents. Unlike the Energy and Voting datasets, where $\beta$ strongly influences compromised discomfort costs, the Privacy dataset shows that scale of adversaries serves as the dominant driver of these costs beyond $\beta>0.3$. This highlights the Privacy dataset heightened sensitivity to adversarial density, requiring more targeted interventions to preserve legitimate agent objectives.

### v2

@@ III-B Network Model
Consider a network with $\mathsf{n}$ agents, denoted by $A$, each identified by a unique ID in the set $\{1,2,3,...,\mathsf{n}\}$. The network topology is represented as a connected graph $G=(A,E)$, where $A$ is the set of agents, and $E$ is the set of edges, with $(j,i)\in E$ not necessarily implying $(i,j)\in E$. Agents interact within a self-organized network through bidirectional communication to exchange information and update their states to align with system goals.
Table III summarizes the notations used throughout the paper to formalize the network model and optimization framework.
TABLE III: Nomenclature utilized in this research
@@ III-C Optimization Framework in Discrete-Choice Scenarios
Agents aim to balance their individual preferences with system-wide goals, which often involve conflicting criteria. Each agent $a$ has an individual preference for its plans, quantified by the discomfort cost ($D$), such that $D$=${}_{a,i}=f_{D}(p_{\mathsf{a,i}})$, where $f_{D}$ measures how undesirable plan $i$ is for agent $a$ based on the agent’s preferences; lower costs indicate more preferred plans. Each agent $a$ evaluates the costs $D$ for each possible plan $p_{\mathsf{a,i}}$$\in$ $P_{\mathsf{a}}$ and the plan with the minimum cost is selected.
While agents aim to minimize their own discomfort, they may also consider system-wide metrics such as the inefficiency cost ($I$). The inefficiency cost ($I$) is the measure used to evaluate the collective system-wide performance based on the aggregated responses of all agents. It represents the system-wide performance inefficiency that agents aim to minimize through coordinated decision-making: $I=f_{I}\left(\sum_{a=1}^{\mathsf{n}}(s_{\mathsf{a}})\right)$. Each agent selects a plan that minimizes a weighted combination of individual discomfort and system-wide inefficiency, as shown in Equation 1.
@@
$=\arg\min_{i=1}^{k}\left[\alpha_{a}\cdot f_{I}\left(s_{1}+s_{2}+\cdots+s_{\mathsf{n}}\right)\right.$
+ $\beta_{a}\cdot f_{D}\left(D_{1,s},D_{2,s},\ldots,D_{\mathsf{n},s}\right)\right]\text{,}$ (1)
@@
The behavior of agent $a$ is modeled by the corresponding weights $\alpha_{a}$ and $\beta_{a}$, which represent agent’s priorities between minimizing system-wide inefficiency and personal discomfort, respectively. A higher weight indicates a higher preference for minimizing the corresponding objective. On the other hand, a weight of 0 means that the corresponding objective is not considered. For instance, an agent with $\alpha_{a}=1$ and $\beta_{a}=0$ behaves altruistically, prioritizing global goals. Conversely, $\alpha_{a}=0$ and $\beta_{a}=1$ define a selfish agent focusing solely on its individual preference.
@@ III-D Adversarial Distributed Optimization Model
We propose an adversarial model applicable across a range of combinatorial optimization scenarios. In this model, the agent population $A$ is partitioned into two disjoint subsets: legitimate agents $A_{\mathsf{l}}$ and adversary agents $A_{\mathsf{d}}$, such that $A=A_{\mathsf{l}}\cup A_{\mathsf{d}}$. While legitimate agents, $A_{\mathsf{l}}\subseteq A$, align their actions with system-wide objectives to optimize overall performance, adversary agents, $A_{\mathsf{d}}\subseteq A$, prioritize their individual interests over collective system goals by adapting their behavior to maximize personal benefits. For instance, in a bike-sharing system [53], optimization ensures a balanced distribution of bikes across stations to meet user demand. Legitimate users may select pick-up and drop-off stations while considering system-wide efficiency, maintaining network equilibrium. In contrast, adversary users prioritize their own convenience, selecting stations solely based on personal preference, leading to imbalances such as empty or overloaded stations, and ultimately degrading overall efficiency and user satisfaction.
Adversarial behavior is modeled by adjusting the weight $\beta_{a}$ in the agent’s decision function (Equation 1). Legitimate agents are assigned $\beta_{l}=0$, fully aligning with system goals, while adversary agents are assigned $\beta_{d}>0$, increasing emphasis on personal discomfort minimization at the expense of system-wide efficiency. This parameterization allows adversarial behavior to be modeled in a continuous space, from fully altruistic to fully selfish.
By varying the distribution and severity of adversarial weights across the agent population, our model enables systematic analysis of resilience, vulnerability, and collapse in distributed optimization. This includes assessing how the impact of adversarial behavior is linked to the network positioning of the adversary agents and legitimate agents. Adversarial behavior amplifies the inefficiency cost $I$, reflecting the trade-off between minimizing individual discomfort ($D$) and optimizing overall system efficiency ($I$). While $D$ focuses on individual preferences, $I$ addresses the system-wide inefficiency caused by deviations from optimal resource allocation, underscoring the conflict between individual and collective optimization goals.
@@ IV Experimental Methodology
This section illustrates the distributed optimization method employed as a case study, the experimental setup, the application scenarios, the measured variables and the evaluation metrics.
@@
IV-A Distributed Optimization Method
The adversarial distributed optimization model is implemented within the Iterative Economic Planning and Optimized Selections (I-EPOS) framework. I-EPOS is a discrete-choice distributed combinatorial optimization algorithm for large-scale multi-agent networks [53, 54]. It employs a self-organized, multi-level hierarchical structure to enable efficient communication, coordination, and scalability while minimizing communication overhead [57].
I-EPOS enables the agents to iteratively coordinate their choices in collective decision-making. Each iteration consists of two distinct phases: a bottom-up phase and a top-down phase. During the bottom-up phase, agents select plans based on the aggregated choices of agents in the branch beneath, as well as the selections made by all agents in the previous iteration. Conversely, the top-down phase addresses incomplete knowledge from higher branches in the hierarchy, enabling agents to revert to previous selections if no cost reduction is achieved. This process continues until a predefined iteration limit is reached or no further improvement in the optimization objective occurs [53, 54]. This iterative coordination mechanism addresses the inherent complexity of multi-agent optimization, particularly under non-linear cost functions and incomplete knowledge of other agents’ choices. These conditions make the optimization problem NP-hard [53], requiring distributed coordination methods that allow agents to refine their decisions based on both local preferences and system-wide impact.
The hierarchical network is structured as an acyclic graph, where each parent agent aggregates responses from its children by avoiding double counting. This design ensures efficient coordination when optimizing individual decisions and system-wide objectives [53, 57, 55]. I-EPOS is well-suited for adversarial scenarios in large-scale distributed systems due to its scalability, adaptability to diverse agent behaviors, and potential mitigate adversarial conditions [54, 58].
To apply the proposed adversarial model, the I-EPOS framework was extended to support heterogeneous agent behaviors. The original implementation assumed uniform agent preferences across the population. We enhanced the framework to allow agents to configure individual decision-making weights. This enhancement enables modeling adversarial agents with varying levels of behavioral severity and is made available as an open-source artifact to support reproducibility and future research [^4] Available at https://github.com/epournaras/EPOS..
@@ IV-B Experimental Setup
Experiments are conducted using multiple HPC servers with varying configurations that support large-scale experimentation and ensure computational efficiency. These include high-memory nodes (up to 768 GB) and multi-core processors (up to 40 cores per node). In addition to these servers, the University of Leeds ARC4 system [^5] ARC4 is an HPC cluster at Leeds providing a Linux-based HPC service based on CentOS 7. More information: https://arcdocs.leeds.ac.uk/systems/arc4.html is utilized. The ARC4 system includes two nodes, each equipped with 40 cores, 768 GB of memory, and 800 GB of storage, providing robust computational capacity for large-scale experiments.
@@ IV-C Application Scenarios
Adversarial distributed optimization is studied in three application scenarios based on real-world data and a synethic dataset. Table IV provides an overview of the datasets used in this research, including the agent populations, the number and size of plans per agent, agent representation and the interpretation of discomfort and inefficiency costs within each application domain. These costs are defined explicitly through the optimization objectives and reflect domain-specific constraints. Further mathematical definitions of cost functions are provided in Appendix A: Cost Function Definitions.
TABLE IV: Description of the datasets and experimental setup in this research
Dataset Name No. Agents No. Plans Plan Size Agents Discomfort Cost Inefficiency Cost Total Experiments Energy 1000 10 144 Households Time shift from intrinsic preference Variance of energy demand 3,118,560 Privacy 72 3 64 Smart phone users Privacy loss Mismatch between shared and desired data 498,780 (2 target signals) Voting 266 31 5 Voters Compromise distance from intrinsic voting preferences Polarization 103,456,800 (120 target signals) Gaussian (synthetic) 10–100 2–10 2 Simulated agents Ranking distance Variance 4,125,000
@@ IV-C2 Privacy dataset
The privacy dataset originates from a living-lab experiment at the Decision Science Laboratory [^7] https://www.descil.ethz.ch of ETH Zurich, involving 72 participants evaluating 64 data-sharing scenarios that involve 4 sensor types, data collectors, and contexts [58]. The data-sharing choices of each participant in the experiment determine three data-sharing plans, representing their intrinsic motivation to share and two rewarded scenarios. The plans are assessed using privacy valuation schemes assigning normalized costs ranging from 0 to 1, where lower costs indicate less privacy compromise. The dataset facilitates testing under high and low privacy-preservation target signals. The inefficiency cost is calculated by the residual sum of squares between the shared and the desired data that measures their mismatch and is an indicator of quality of service supported by the collected data. Adversarial participants disrupt coordination by focusing solely on minimizing their data sharing, under-mining the quality of service of data collectors.
@@ IV-C3 Voting dataset
This new dataset is derived from voting data in a regional election with five candidates and 266 voters [^8] The UK Labour Party Leadership Vote Available at https://preflib.simonrey.fr/datasets [59]. Each voter has 31 alternative voting plans, representing ranked preferences among the five candidates. The optimization focuses on minimizing polarization in the voting outcomes, which refers to reaching the same voting outcome but with compromises that reduce polarization. Polarization here is the inefficiency cost and it refers to the mismatch from a liner ranking of the alternatives in the voting outcomes, although other polarization models could be studied as well [60]. The rationale of linearity is to deviate from concentrating the voters’ preferences to two opposing poles. To control for the same voting outcome, 120 target signals are generated from all combinations of values 0, 0.25, 0.5, 0.75, and 1, representing the linear ranking of alternatives. Adversarial behavior occurs when voters prioritize their intrinsic preferences, i.e., no compromises to reduce polarization.
IV-C4 Synthetic Gaussian dataset
The synthetic dataset is constructed to evaluate system scalability under controlled, domain-agnostic conditions. It includes 100 agents with 10 generated plans, where each plan is a 100-dimensional vector sampled from a Gaussian distribution $\mathcal{N}(0,1)$. Plans are sorted by their index, with lower indices arbitrary indicating higher agent preference. Discomfort cost is defined as the rank of the selected plan—i.e., a higher index reflects greater deviation from the most preferred option. Inefficiency cost is measured as the variance of the aggregated global response, capturing system-wide imbalance. This synthetic setup allows systematic analysis of adversarial effects across varying agent populations, number/size of plans, and attack configurations.
@@ IV-D Varying Dimensions and Performed Experiments
The following dimensions are studied in the performed experiments:
@@
  • Adversarial severity ($\beta$): The adversarial preference to minimize discomfort cost ($\beta$) is varied across 30 levels, with $\beta$ incrementing from 0 to 1 such that $\beta=\frac{b}{30}$ for $b=\{1,2,3,...,30\}$.
  • Adversary position: The influence of adversary positions within the hierarchical network is analyzed using two approaches: layer-wise and cumulative structural analysis. A binary tree structure is employed, with each hierarchical layer containing approximately $\log_{2}|A|$ agents, where $|A|$ is the total number of agents. The structural analysis evaluates inefficiency costs under varying adversary scales (25%, 50%, 75%, 100%) at each layer of the hierarchy. The cumulative analysis examines the aggregated impact of adversary agents positioned incrementally in top-down (root-to-leaf) and bottom-up (leaf-to-root) configurations.
For each dataset, experiments are conducted across 30 adversarial severity levels ($\beta$) with 100 simulation runs per configuration.
Layer-wise structural analysis is performed at four adversarial proportions $p\in\{25\%,50\%,75\%,100\%\}$ within each layer of the hierarchical topology, where the number of layers is defined as $\lceil\log_{2}|A|\rceil$. For each layer $L$, the number of adversarial agents is calculated as $k_{p}=\max\left(1,\ \left\lceil\frac{p}{100}\cdot|A_{L}|\right\rceil\right)$, where $|A_{L}|$ denotes the number of agents in layer $L$. Adversarial configurations are sampled up to 100 combinations per setting to ensure computational feasibility [^9] The equation assumes up to 100 combinations per layer-percentage, but the binary hierarchy often yields fewer due to limited agents per layer.. In addition, two cumulative structural attack scenarios are simulated, top-down (root-to-leaf) and bottom-up (leaf-to-root), introducing $2\times|A|$ further experiments per dataset. The total number of experiments per dataset is calculated as:
Total Experiments=(30×number of $\text{Total Experiments}=(30\times\text{number of signals})\times\left[(100\times|A|)+\sum_{L}\sum_{p}\min\left(100,\ \binom{|A_{L}|}{k_{p}}\right)+(2\times|A|)\right]$
For the synthetic Gaussian dataset, experiments vary both agent populations and the number of plans per agent (from 2 to 10). The total number of experiments is computed as: $\text{Total Experiments}_{\text{Gaussian}}=\sum_{i=1}^{10}\left(10i\times 30\times 5\times 50\right),$ where $i$ is the number of agents, 30 is the number of severity levels, 5 is the number of plans, and 50 is the number of random structural permutations (i.e., reordering of agents in the tree).
@@ IV-E Evaluation Metrics
Optimization objectives: System performance is assessed using three metrics: inefficiency cost $I$, discomfort cost $D$, and the compromise cost of legitimate agents. Inefficiency cost captures the deviation from optimal system performance. Discomfort cost reflects individual agents’ dissatisfaction, i.e. to what extent a plan is not the most preferred one. The compromise cost quantifies the increase in discomfort experienced by legitimate agents due to adversarial influence. It is calculated as the difference in discomfort between scenarios with and without adversaries, highlighting the collective burden legitimate agents bear to maintain system performance.
Pareto optimality: In multi-objective optimization, the Pareto front defines solutions where no objective can improve without compromising another objective. The knee point on this front identifies the most balanced trade-off between competing objectives. This study uses the Minimum Manhattan Distance (MMD) method to locate the knee point, measuring the distance from each Pareto solution to an ideal reference point where both objectives are optimized. The solution with the smallest distance is selected. This approach ensures a balanced trade-off between discomfort ($D$) and inefficiency ($I$) in line with established approaches [61, 62].
Resilience, vulnerability, and collapse framework: To classify system states, the multi-Otsu thresholding method is applied to segment inefficiency and discomfort values into three distinct regions: resilience (low inefficiency), vulnerability (moderate inefficiency), and collapse (high inefficiency). This technique minimizes intra-class variance, offering a robust framework to detect transitions in system performance under adversarial conditions [63, 64].
@@
(c) Compromised discomfort costs for legitimate agents
Figure 1: Inefficiency, discomfort, and compromised discomfort over adversary scales and severity in the energy, voting, and privacy datasets, , including Pareto knee points and resilience (R), vulnerability (V), and collapse (C) thresholds.
V Results Analysis and Discussion
This section presents the results of extensive experiments evaluating the impact of adversarial agents on multi-objective distributed optimization across the real‐world (energy, voting, and privacy) and synthetic datasets.